# Capsule Network Projectors are Equivariant and Invariant Learners

### v2

@@ B.1 CapsIE Pre-training
Our proposed CapIE model is comprised of a ResNet-18 encoder, SR-CapsNet comprised of a primary capsule layer routed to a second capsule layer. The SR-CapsNet projector takes as input the activation map output of the ResNet prior to the final global average pooling. The predictor network employed is that described in [14]. All details of the architectural design are given in the main paper.
@@ B.2 Downstream Evaluation
B.3 Quantitative Equivariant Evaluation Implementation Details
We evaluate the equivariant properties of the predictor in line with that proposed in [14] reporting the Mean Reciprocal Rank (MRR) and Hit Rate at k (H@k) on the multi-object setting. Given a source and target pose of an object, we first compute the embeddings of each image, and pass the source embedding through the predictor and use the resulting vector to retrieve the nearest neighbours.
The Prediction Retrieval Error (PRE) gives an evaluation of predictor quality, and is given by the distance between its rotation $q_{1}\in\mathbb{H}$ and the target rotation $q_{2}$ as $d=1-<q_{1},q_{2}>^{2}$ of the nearest neighbour of the predicted embedding averaged over the whole dataset. Full implementation details can be found in the open-source code provided.
B.4 Supervised Training of SR-Caps
@@
Appendix C Further Experimentation
C.1 Capsule Downstream Classification on Embeddings.
Table 8 provides the downstream classification evaluation for the frozen output embeddings of the CapsNet projection head. The drop in performance is an expected result inline with that demonstrated by [7, 4]. These results signify the importance of the projection head and specifically its role in decorrelating the embeddings from representations to avoid overfitting to the pre-training objective. We do note however, that our observed performance drop is inline or slightly less than that reported in [14].
Table 8: Capsule downstream invariance evaluation on projection head embeddings.
Embedding Dims Classification (Top-1%)
Method Inv. Equi. Representations Embeddings
Supervised - - 87.47
SIE 512 512 82.94
Capsule Projector Naive Evaluation
CapsIE - 16 16 256 78.96 65.83
CapsIE - 32 32 512 80.00 69.12
CapsIE - 64 64 1024 80.26 56.64
C.2 Invariant and Equivariant SSL Benchmarks
We report below the classification (invariant, Table 9), rotation prediction, and colour prediction (equivariant, Table 10) performance of baseline self-supervised methods. The below baseline results are acquired from [14], with exception to those denoted by ‘*’ which corresponds to our re-implementation.
@@ C.2 Invariant and Equivariant SSL Benchmarks
CapsIE - 16 16 256 78.96 - -
CapsIE - 32 32 512 80.00 - -
CapsIE - 64 64 1024 80.26 - -
@@
CapsIE - 16 16 256 - 82.96 -
CapsIE - 32 32 512 - 83.49 -
CapsIE - 64 64 1024 - 83.64 -
@@
CapsIE - 16 0.78 - - 0.97 - -
CapsIE - 32 0.75 - - 0.97 - -
CapsIE - 64 0.72 - - 0.97 - -
@@
CapsIE - 16 - - 0.78 - - 0.97
CapsIE - 32 - - 0.77 - - 0.97
CapsIE - 64 - - 0.78 - - 0.97

### v3

@@ B.1 CapsIE Pre-training
Our proposed CapIE model is comprised of a ResNet-18 encoder, SR-CapsNet comprised of a primary capsule layer routed to a second capsule layer. The SR-CapsNet projector takes as input the activation map output of the ResNet prior to the final global average pooling. The predictor network employed is that described in (Garrido et al., 2023). All details of the architectural design are given in the main paper.
@@ B.2 Downstream Evaluation
B.3 Supervised Training of SR-Caps
@@
B.4 Invariant and Equivariant SSL Benchmarks
We report below the classification (invariant, Table 7), rotation prediction, and colour prediction (equivariant, Table 8) performance of baseline self-supervised methods. The below baseline results are acquired from (Garrido et al., 2023), with the exception of those denoted by ‘*’ which corresponds to our re-implementation.
@@ B.4 Invariant and Equivariant SSL Benchmarks
CapsIE - 16 16 256 76.51 - -
CapsIE - 32 32 512 79.14 - -
CapsIE - 64 64 1024 79.60 - -
@@
SIE 1024 1024 - 80.53 77.64
CapsIE - 16 16 256 - 74.90 -
CapsIE - 32 32 512 - 78.60 -
CapsIE - 64 64 1024 - 79.24 -
@@
CapsIE - 16 0.68 - - 0.02 - -
CapsIE - 32 0.74 - - 0.04 - -
CapsIE - 64 0.72 - - 0.01 - -
@@
SIE - 0.38 0.58 - 0.45 0.09
CapsIE - 16 - - 0.64 - - -0.01
CapsIE - 32 - - 0.71 - - -0.04
CapsIE - 64 - - 0.67 - - -0.09